the mean weight of baggage of 25 passengers from usa bound for the philippines is 31.8 kilos with a variance of 16. test the claim that the population mean baggage weight is equal to 30 kilos at 0.02 level of significance.


1
Expert's answer
2021-09-28T15:41:26-0400

Null hypothesis: "u=30"

Alternative hypothesis: "u\not =30"

"\alpha = 0.02"

Test statistic:


"Z = {\frac {(x - u)*\sqrt{n}} {s}}"

where x - sample mean, s - sample standart deviation(square root of the variance) , n - sample size.

In our case "Z={\frac {(31.8-30)*5} {4}}=" 2.25

p-value: "p=2P(Z>2.25)=2*0.01222=0.02444"

p-value is greater than significance level, so there is no evidence to reject the null hypothesis that the population mean baggage weight is equal to 30.